Can you let me know if I answered this math problem correctly?
If Sin theta = -2/3 find Sin of (-theta)
I came up with Sin theta = 2/3.
Thank you.
Sarah McGuire

Found 3 solutions by Nate, stanbon, funmath:
Answer by Nate(3500)

Define:
f(theta) = sin(theta) = -2/3
f(-theta) = sin(-theta) = -sin(theta) = -(-2/3) = 2/3

Answer by stanbon(48564)

If Sin theta = -2/3 find Sin of (-theta)
----------
Formula: sin(x)=-sin(-x)
Therefore If sin theta = -2/3,
sin(-theta)= - sin theta = 2/3
Cheers,
Stan H.

Answer by funmath(2873)

Hi Sarah,
sin theta is an odd, so you're right,
if sin(theta)=-2/3, then sin(-theta)=2/3
Keep up the good work!
Happy Calculating!!!